OpenStudy (anonymous):

I need a step by step proving of the statement If c is any nonzero real number then c/c = 1

OpenStudy (anonymous):

Starting from...?

OpenStudy (anonymous):

i guess it depends of what you definition of division is. if you say \[\frac{a}{b}=c \iff a=bc\] then \[\frac{c}{c}=x \iff xc = c \iff x = 1\]

OpenStudy (anonymous):

you have to start with something, as estudier said

OpenStudy (anonymous):

I don't know where to start T_T

OpenStudy (anonymous):

Well, satellite73 has linked together the operations of division and multiplication to conclude that c/c = 1. Does his argument not convince you?

OpenStudy (anonymous):

it is totally obvious that any non-zero number divided by itself is 1. when asked this question, are you given a definition for division?

OpenStudy (anonymous):

Yes, I was.

OpenStudy (anonymous):

How about c * 1/c? Maybe we could do something with that?

OpenStudy (anonymous):

yes,yes ...I remember, our instructor told us about Multiplicative Inverses

OpenStudy (anonymous):

may stuff is ending up all over the place

OpenStudy (anonymous):

Now we're talking...reckon you can do it now...

OpenStudy (anonymous):

Yey..thank you so much for the help :) Satellite and estudier

OpenStudy (anonymous):

you are welcome, what is this course you are doing?

OpenStudy (anonymous):

It's College Algebra and Trigonometry

OpenStudy (anonymous):

Ah, right...Good luck with it then.

OpenStudy (anonymous):

Thanks! :)